OpenStudy (anonymous):

log x + log(x-3)=1 help!

OpenStudy (anonymous):

\[\log x(x-3)=1\]

OpenStudy (anonymous):

\[\log a + \log b = \log ab\]\[\log a - \log b = \log a/b\]

OpenStudy (campbell_st):

simplify the logs 1st \[\log(a) + \log(b) = \log(ab)\] you need to simplify the left hand side using the law. next raise both sides to the power of 10. Assuming that its a base 10 log. the law says \[10^{\log_{10} (a)} = a\] and the right hand side will be 10^1 now solve for x

OpenStudy (anonymous):

\[10^1=x(x-3)\]\[10=x^2-3x\]\[0=x^2-3x-10\]\[0=(x-5)(x+2)\]\[x=5,-2\]

OpenStudy (mathstudent55):

log [x(x - 3)] = 1 log(x^2 - 3x) = 1 Using def of log: x^2 - 3x = 10^1 x^2 - 3x - 10 = 0 (x - 5)(x + 2) = 0 x = 5 or x = -2 Since log(-2) is not defined, x = 5

OpenStudy (anonymous):

\[\log_{10} x=y \rightarrow 10^y=x\]

OpenStudy (campbell_st):

you need to ignore -2 as a solution as you can't take the log of a negative number.

OpenStudy (anonymous):

Good point

OpenStudy (anonymous):

So basically when you get rid of the logs from one side, the other side becomes 10?

OpenStudy (campbell_st):

yes... or the base of the logs is it was base e..... e.g ln(x) = 5 x = e^5

OpenStudy (anonymous):

ok got it thanks!

OpenStudy (anonymous):

if not written otherwise. log is always base 10. :) Good luck

OpenStudy (anonymous):

thank you!